# What Makes a Visualization Complex?

### v1

@@ Abstract
We investigate the perceived visual complexity (VC) in data visualizations using objective image-based metrics. We collected VC scores through a large-scale crowdsourcing experiment involving 349 participants and 1,800 visualization images. We then examined how these scores align with 12 image-based metrics spanning information-theoretic, clutter, color, and our two object-based metrics. Our results show that both low-level image properties and the high-level elements affect perceived VC in visualization images; The number of corners and distinct colors are robust metrics across visualizations. Second, feature congestion, an information-theoretic metric capturing statistical patterns in color and texture, is the strongest predictor of perceived complexity in visualizations rich in the same stimuli; edge density effectively explains VC in node-link diagrams. Additionally, we observe a bell-curve effect for text annotations: increasing text-to-ink ratio (TiR) initially reduces complexity, reaching an optimal point, beyond which further text increases perceived complexity. Our quantification pipeline is also interpretable—enabling metric-based explanations— grounded in the VisComplexity2K dataset, bridging computational metrics with human perceptual responses. osf.io/5xe8a has the preregistration and osf.io/bdet6 has the VisComplexity2K dataset, source code, and all Apdx. and figures.
@@ Introduction
Evaluating perceived VC can also contribute to foundational research in visualization. VC has been defined as “the amount of detail or intricacy” [129, 107]. Most evaluation use simple images, that are less complex than those found in practical, real-world uses [156]. In contrast, practitioners foraging information need to handle in realistic visual design to achieve more generalizable outcomes [150]. This gap introduces at least two evaluation challenges. First, the vast majority of visualization images contain an array of features, patterns, and contexts. We have a limited understanding of their feature space. Even though high-level features (e. g., number of elements [86], text [49], and the scene gist [99]) and a computational equivalent (e. g., feature congestion [118]) contribute to VC, studied previously primarily for natural scenes [130], it is unclear if any of the existing metrics and theories are generalizable across domains. Second, we also lack of understanding of observers’ experiences when examining these images.
@@
Our goal is to systematically identify some of the image features that influence VC. In this work, we explore how VC can be viewed through both perceived and objective lenses. To this end, we draw inspiration from the extensive works in vision science that collects large-scale absolute VC scores [124]. Furthermore, we draw on a precedent for such metric-based explanations [107], to study what metrics can represent VC judgment. In the longer term, measuring and explaining human judgment with an extensive set of metrics can reverse engineer high-level multifaceted understanding of the underlying mechanisms [72].
@@
Method (Figure 1), we begin with an extensive literature review on subjective VC factors in Table 1 to guide our selection of 12 objective metrics in Table 2. Some metrics, such as information-theoretic metrics, capture mathematical uncertainty, while others are sensitive to clutter (feature congestion) [120], shape and structure, or color. We further introduce two new metrics to represent visualization elements: (1) MeC (the number of meaningful colors), to represent the number of component parts in an image—by applying color-as-object-boundaries [41]; (2) TiR (text-ink-ratio)—the proportion of annotation in a visualization image, to investigate how context influences VC, given its prominent role in visualization [109] and its power to assist viewers to comprehend concepts that are otherwise difficult to interpret. We collect perceived VC scores of 1,800 visualization images and then use image-based feature metrics to capture VC judgment. While a precedent for metric-based explanations of specific visualizations (e. g., networks [107]) and stimulus-specific clutter (e. g., color and texture [120]) exists within our community, measuring and explaining human perceived VC using an extensive set of metrics offers a pathway towards finding a set of low-dimensional proxy to engineer high-dimensional VC.
@@
Results. We collected a large set of perceived VC scores from 1,800 images (What Makes a Visualization Complex?). Our analysis reveals that a range of metrics contribute to explaining perceived VC. Among them, edge density (ED) emerges as a dominant factor across visualization images. Object-level metrics such as color count (MeC) and text-ink-ratio (TiR) capture high-level aspects of visual complexity but have little correlation with pixel-based metrics. A notable finding is that clutter-based feature congestion (FC) [120, 118], an information-theoretic measure adapted to account for statistical scene structures of color and texture, proves particularly effective in distinguishing VC in visualizations rich in the same color and texture variations. Finally, we observe that the amount of annotation text, as measured by TiR, has a non-linear U-shaped relation with perceived VC: moderate levels of text reduce VC while excessive annotations increase it, extending the current literature.
@@
In summary, our main contributions are: (1) A crowdsourcing study collecting perceived VC scores for 1,800 visualization images from 349 participants. (2) A metrics study that analyzes how 10 existing and 2 new metrics help us interpret perceived VC. (3) We provide the human and metric scores from both studies, along with the 1,800 visualization images, in a dataset called VisComplexity2K, openly available at osf.io/bdet6.
@@ I Background and Related Work
This section reviews the definitions of VC and the associated metrics. As outlined in Table 1, our literature review identifies five key categories, which serve as the foundation for organizing the metrics summarized in Table 2, largely from vision science.
@@ I.1 Definitions of Visual Complexity
| Information theoretic | Information amount in a mathematical sense [127]. | [92, 99, 124, 22, 65, 67, 150] |
| Clutter | There exists “little room for a new feature to draw attention”[120]. | [77, 124, 120, 99] |
| Color | The variety of colors (hue, saturation, and brightness), the level of continuity, intensity, and range of colors. | [59, 107, 120, 92, 27, 87, 38] |
| Shape and Structure | The variety and intricacy of edges and turns within an image, encompassing the contours and boundaries and Structural clarity (e.g., regularity, heterogeneity, region segmentation by color or textures). | [87, 92, 5, 107, 77, 132, 27, 100, 110, 59] |
@@
| Info.-theoretic (O.Info) | (1) | Shannon Image entropy (O.IE) | The amount of information or randomness in the image, calculated by applying the entropy formula to the grayscale pixel intensity values of the image. | , where , the probability of occurrence of the grayscale pixel intensity in the image, and , the total number of different pixel intensities. | [40] |
| (2) | Kolmogorov Complexity (O.KC) | Image size after removing redundant information. | , where , the original size of the image, and , the size of redundant compressible information. | [25] | |
| (3) | Subband entropy (O.SE) | The entropy of the image content within subbands, calculated using the probability distribution of the the wavelet coefficients within each subband. | , where , the probability distribution of intensity values or coefficients within the subband, and the sum runs over all subbands. | [120, 119] | |
| (4) | Information gain (O.IG) | The amount of additional information needed based on the mean information gain to determine the color of an adjacent cell given the color of a particular cell in a 2D pattern. | , where , the joint probability that a given cell has the color and the adjacent cell has the color, and , the conditional probability that the adjacent cell has the color given that a cell has the color. | [3] | |
| Clutter (O.CL) | (5) | Feature congestion (O.FC) | The level of clutter in an image by calculating color, luminance contrast, and orientation. | , where , the proportion of energy of the feature relative to the total energy of all features in the image. | [120] |
| (6) | Heterogeneity (O.H) | Calculated from Gray-Level Concurrent Matrix to provide insights into the texture and contrast of an image. Higher heterogeneity values indicate more contrast and variation in gray levels, suggesting a more complex or textured image. | , represents the probability that a pixel with intensity occurs adjacent to a pixel with intensity , and is the total number of gray levels in the image. | [29] | |
@@
| (8) | Color RGB entropy (O.ERGB) | The variability in the distribution of colors calculated by applying Shannon’s entropy to the color histogram of an image. | , where , the probability of occurrence of the color in the histogram, and , the total number of different colors histogram. | [40] | |
@@
| (12) | Number-of-color (O.MeC) | The number of meaningful colors count. | The set of human namable colors of Heer and Stone [50] in an image. |
@@ I.2 Five Factors that Govern Visual Complexity
I.2.1 Data-driven Information Theoretic Measure
The information-theoretic measure, often computed using Shannon’s entropy [127], is a purely mathematical representation of pixel data in images. In data visualization, Yang-Peläez and Flowers (Y-F) [155] extended Shannon’s classic entropy by partitioning ‘information’ into three types: syntactic (marks, lines, regions, and structures thereof), semantics (meaning of those marks or other syntactic information), and pragmatic (the value or usefulness of that information). The number of bits is subsequently derived from marks by searching neighboring pixels. Every bit/pixel carries an equal weight representing the meaning of the message. Later, Chen and Jänicke [22] (C-J) conceptualize the Y-F method to take data variations in every stage of the visualization workflow, by weighting the neighboring pixels to compute the information capacity from data in an image. Both Y-F and C-J’s methods require knowing the information flow in the image space. To automate the process to compute data-bits, Jänicke et al. [67, 80] use color entropy to highlight and sort flow features. Here, they treated sensory perception independent of data saliency to analyze flow features by information theoretic solutions. Furthermore, Jänicke and Scheuermann [66] computed statistical complexity in data [67] to compare, analyze, and highlight features that assist human observers [91].
What inspired us in these works is the transition from the low-level design element to key probabilistic data-specific knowledge in images, modeled under the same information theoretic framework of mathematic uncertainty. However, we were unable to directly extend their work, as there is currently no established method for defining bits in the context of general visualizations beyond flow fields. A broader idea drawn from this line of research is that more complex images tend to exhibit lower redundancy and thus higher entropy. For this reason, the compression rate of an image—aligned with Kolmogorov complexity [115]—serves as reasonable approximation of pixel-level uncertainty (Item (1) O.KC, Table 2). Building on this idea, We introduced several additional metrics (Items (2)-(4), Table 2) that adopt this entropy-based approach. Furthermore, we incorporated pixel-level RGB entropy into our metric set (Item (8) O.ERGB, Table 2).
@@ I.2.2 Clutter-based Feature Congestion
The design of FC originated from an interesting question: whether the robust measure of the “number of unique objects” [17, 86] can be computed mathematically [118] to eliminate the need for labor-intensive manual coding. Here, congestion is characterized by the lack of available space for new features [120], and “the more cluttered a display or scene is, the more difficult it would be to add a new item that would reliably draw attention” [118]. Subsequently, this metric assembles the spatial frequency of Gabor texture patches and colors to simulate receptive fields, extending the subband entropy onto the object-level structure of color, orientation, and saliency [118]. This is grounded on the understanding that an image is considered ‘simple’ [99] when it is symmetric and structured. Thus, complexity is inversely correlated with perceptual grouping and regularity [53]. Empirical evaluation results also supported the FC metric (Item (5) O.FC, Table 2), subband entropy (Item (3) O.SE, Table 2) and edge density (Item (9) O.ED, Table 2) on measuring VC [118, 117]. Borgo et al. [9] further emphasized the idea of statistical attributes to characterize complexity. This line of work informs us that information amount influences VC. Both Rosenholtz in vision science and Jänicke and Chen in visualization use color as the primary channel for information-theoretic metrics [117, 67]. Given that color and textures are uniquely represented in visualization images, we modeled between continuity and complexity conditioned for continuous and discrete color and texture stimuli, or RR-ColorTexture of Continuous Heatmap and Discrete Heatmaps in images.
@@ I.2.3 Color-based Measures
Color also plays a significant role in shaping perceptual experiences. For example, Li and Chen [79] suggest that using too many colors can reduce memorability due to increased clutter.
Moreover, color is a leading factor in how visual groupings are perceived, supporting both segmentation and the identification of structure within complex visual scenes. In vision science, both Greene et al. [41] and Rosenholtz [117] highlight the role of color in defining object boundaries. Many studies have linked color-related features to VC. For example, Rosenholtz et al.observed that, in natural object scenes, color variability is independent of both edge density and subband entropy [118]. Both Oliva et al. [99] and Rosenholtz et al. [120] reported that perceived VC increases more with color variability (the number of colors and how different they are) than with the presence of high frequency details in color channels. Corchs [27, 55] introduced metrics such as color harmony and color coherence to show that both the interaction and uniformity of colors contribute to VC. Here, we balance our metric choices to capture both statistical color properties and object-level color counts. Specifically, we used colorfulness (Item (7) O.CF, Table 2), defined as “the variety and diversity of colors present in the region” [46], and gray-level color heterogeneity (Item (6) O.H, Table 2) to indicate overall saturation [46], in addition to the aforementioned O.ERGB.
I.2.4 Object-based Measure
Several studies have focused on object-level measures, emphasizing that VC involves higher-level cognitive processing, which requires the analysis and integration of multiple elements along with their semantic content. Semantic contents in visualization can have two types: graphical elements (e. g., legend, tick marks); and annotations [109]) conveying key insights and takeaways [1]. The effect of texts on VC can be mixed. For one, they can effectively guide viewing behavior [14, 11] to reduce VC. For the other and from a pixel-based perspective, text is treated as a visual structure—often dark-colored, densely arranged, and high in contrast, that can increase complexity. (Item (11) O.TiR, Table 2.)

### v3

# What Makes a Visualization Image Complex?

@@ Abstract
We investigate the perceived visual complexity (VC) in data visualizations using objective image-based metrics. We collected VC scores through a large-scale crowdsourcing experiment involving 349 participants and 1,800 visualization images. We then examined how these scores align with 12 image-based metrics spanning pixel-based and statistic-information-theoretic (clutter), color, shape, and our two new object-based metrics (meaningful-color-count (MeC) and text-to-ink ratio (TiR)). Our results show that both low-level edges and high-level elements affect perceived VC in visualization images; the number of corners and distinct colors are robust metrics across visualizations. Second, feature congestion, a statistical information-theoretic metric capturing color and texture patterns, is the strongest predictor of perceived complexity in visualizations rich in the same continuous color/texture stimuli; edge density effectively explains VC in node-link diagrams. Additionally, we observe a bell-curve effect for texts: increasing TiR initially reduces complexity, reaching an optimal point, beyond which further text increases VC. Our quantification model is also interpretable—enabling metric-based explanations—grounded in the VisComplexity2K dataset, bridging computational metrics with human perceptual responses. osf.io/5xe8a has the preregistration and osf.io/bdet6 has the VisComplexity2K dataset, source code, and all Apdx. and figures.
@@ Introduction
Evaluating perceived VC can also contribute to foundational research in visualization. VC has been defined as “the amount of detail or intricacy” [130, 107]. Most evaluations use simple images, that are less complex than those found in practical, real-world uses [158]. In contrast, practitioners foraging for information need to handle realistic visual designs to achieve more generalizable outcomes [152]. This gap introduces at least two evaluation challenges. First, the vast majority of visualization images contain an array of features, patterns, and contexts. We have a limited understanding of their feature space. Even though high-level features (e. g., number of elements [85], text [47], and the scene gist [98]) and a computational equivalent (e. g., feature congestion [119]) contribute to VC, studied previously primarily for natural scenes [131], it is unclear if any of the existing metrics and theories are generalizable across domains. Second, we also lack an understanding of observers’ experiences when examining these images.
@@
Our goal is to identify some of the image features that influence VC. In this work, we explore how VC can be viewed through both perceived and objective lenses. To this end, we draw inspiration from the extensive work in vision science that collects large-scale absolute VC scores [124]. Furthermore, we draw on a precedent for such metric-based explanations [107, 117, 110], to study what metrics can represent VC. In the longer term, measuring and explaining a human judgment with an extensive set of metrics can help reverse engineer high-level multifaceted understanding of the perceptual mechanisms [70].
@@
Method (Figure 1). We begin with an extensive literature review on subjective VC factors in Table 1 to guide our selection of 12 objective metrics in Table 2. Some metrics, such as information-theoretic metrics, capture mathematical uncertainty, while others are sensitive to clutter (feature congestion) [120], shape and structure, or color. We further introduce two new metrics to represent visualization elements: (1) MeC (the number of meaningful colors), to represent the number of component parts in an image—by applying color-as-object-boundaries [39]; (2) TiR (text-ink-ratio)—the proportion of annotation in a visualization image, to investigate how context influences VC, given its prominent role in visualization [109] and its power to assist viewers to comprehend concepts that are otherwise difficult to interpret. We collect perceived VC scores of 1,800 visualization images and then use image-based feature metrics to capture VC judgment. While a precedent for metric-based explanations of specific visualizations (e. g., networks [107]) and stimulus-specific clutter (e. g., color and texture [119]) exists within our community, measuring and explaining human perceived VC using an extensive set of metrics offers a pathway towards finding a low-dimensional proxy to engineer high-dimensional VC.
@@
Results. We collected perceived VC scores from 1,800 images ( What Makes a Visualization Image Complex?). Our analysis reveals that a range of metrics contribute to explaining perceived VC. Among them, edge density (ED) emerges as a dominant factor across visualization images. Object-level metrics such as color count (MeC) and text-ink-ratio (TiR) capture high-level aspects of visual complexity but have little correlation with pixel-based metrics. A notable finding is that clutter-based feature congestion (FC) [120, 119], an information-theoretic measure adapted to account for statistical scene structures of color and texture, was particularly effective in distinguishing VC in visualizations rich in the same color and texture variations. Finally, we observe that the amount of annotation text, as measured by TiR, has a non-linear U-shaped relation with perceived VC: moderate levels of text reduce VC while excessive annotations increase it, extending the current literature.
@@
In summary, our main contributions are: (1) A crowdsourcing study collecting perceived VC scores for 1,800 visualization images from 349 participants. (2) A metrics study that analyzes how 10 existing and two new metrics help us interpret perceived VC. (3) We provide the human and metric scores from both studies, along with the 1,800 visualization images, in a dataset called VisComplexity2K, openly available at osf.io/bdet6.
@@ I Background and Related Work
This section reviews the definitions of VC and the associated metrics. As outlined in Table 1, our literature review identifies five key categories, which serve as the foundation for organizing the metrics summarized in Table 2.
@@ I.1 Definitions of Visual Complexity
| Info.-theoretic | Information amount as the size of digital file, inverse of [128], and as saliency. | [92, 98, 124, 22, 63, 65] |
| Clutter (Info.-theoretic) | Summary statistics over the entire image, e. g., there exists “little room for a new feature to draw attention” [120]. | [124, 120, 98] |
| Color | The variety of colors (hue, saturation, and brightness), the level of continuity, intensity, and range of colors. | [107, 120, 37] |
| Shape and Structure | The variety and intricacy of edges and turns within an image, encompassing the contours and boundaries and Structural clarity (e. g., regularity, heterogeneity, region segmentation by color or textures). | [5, 107, 75, 134, 27, 99, 110] |
@@
| Info.-theoretic (O.Info) | (1) | Shannon Image entropy (O.IE) | The amount of information or randomness in the image, calculated by applying the entropy formula to the grayscale pixel intensity values of the image. | , where , the probability of occurrence of the grayscale pixel intensity in the image, and , the total number of different pixel intensities. | [127, 136] |
| (2) | Kolmogorov complexity (O.KC) | Image size after removing redundant information. | , where , the original size of the image, and , the size of redundant compressible information. | [25, 105] | |
| (3) | Subband entropy (O.SE) | The entropy of the image content within subbands, calculated using the probability distribution of the wavelet coefficients within each subband. | , where , the probability distribution of intensity values or coefficients within the subband, and the sum runs over all subbands. | [120] | |
| (4) | Information gain (O.IG) | The amount of additional information needed based on the mean information gain to determine the color of an adjacent cell given the color of a particular cell. | , where , the joint probability that a given cell has the color and the adjacent cell has the color, and , the conditional probability that the adjacent cell has the color given that a cell has the color. | [3] | |
| Clutter (O.CL) | (5) | Feature congestion (O.FC) | The clutter estimate in the image computed by assessing feature viability based on the entropy of color, luminance contrast, and orientation. | , where , the proportion of energy of the feature relative to the total energy of all features in the image. | [120] |
| (6) | Heterogeneity (O.H) | A gray-level concurrent matrix to represent the texture and contrast of an image. Higher heterogeneity values indicate more contrast and variation in gray levels, suggesting a more complex or textured image. | , represents the probability that a pixel with intensity occurs adjacent to a pixel with intensity , and is the total number of gray levels in the image. | [29] | |
@@
| (8) | Color RGB entropy (O.ERGB) | The variability in the distribution of colors calculated by applying Shannon’s entropy to the color histogram of an image. | , where , the probability of occurrence of the color in the histogram, and , the total number of different colors histogram. | [127, 136] | |
@@
| (12) | Number-of-color (O.MeC) | The number of meaningful colors count. | The curated human namable colors of Heer and Stone [48] in an image. |
@@ I.2 Five Factors that Govern Visual Complexity
I.2.1 Data-driven Information-Theoretic Measure
The information-theoretic measure, often computed using Shannon’s entropy [128], is a purely mathematical representation of pixel data in images. In data visualization, Yang-Peläez and Flowers (Y-F) [157] extended Shannon’s classic entropy by partitioning ‘information’ into three types: syntactic (marks, lines, regions, and structures thereof), semantics (meaning of those marks or other syntactic information), and pragmatic (the value or usefulness of that information). The number of bits is subsequently derived from marks by searching neighboring pixels. Every bit/pixel carries an equal weight representing the meaning of the message. Later, Chen and Jänicke [22] (C-J) conceptualized the Y-F method to take data variations in every stage of the visualization workflow, by weighting the neighboring pixels to compute the information capacity from data in an image. Both Y-F and C-J’s methods require knowing the information flow in the image space. To automate the process to compute data-bits, Jänicke et al. [65, 79] used color entropy to highlight and sort flow features. Here, they treated sensory perception independent of data saliency to analyze flow features by information-theoretic solutions. Furthermore, Jänicke and Scheuermann [64] computed statistical complexity in data [65] to compare, analyze, and highlight features that assist human observers [90].
What inspired us in these works is the transition from the low-level design element to key probabilistic data-specific knowledge in images, modeled under the same information-theoretic framework of mathematical uncertainty. However, we were unable to directly extend their work, as there is currently no established method for defining bits in the context of general visualizations beyond flow fields. A broader idea drawn from this line of research is that more complex images tend to exhibit lower redundancy and thus higher entropy. For this reason, the compression rate of an image—aligned with Kolmogorov complexity [115]—serves as reasonable approximation of pixel-level uncertainty (Item (1) O.KC, Table 2). Building on this idea, we introduced several additional metrics (Items (2)-(4), Table 2) that adopt this entropy-based approach. Furthermore, we incorporated pixel-level RGB entropy into our metric set (Item (8) O.ERGB, Table 2).
@@ I.2.2 Clutter-based Feature Congestion
The design of FC originated from an interesting question: whether the robust measure of the “number of unique objects” [16, 85] can be computed mathematically [119] to eliminate the need for labor-intensive manual coding. Here, congestion is characterized by the lack of available space for new features [120], and “the more cluttered a display or scene is, the more difficult it would be to add a new item that would reliably draw attention” [119]. Subsequently, this metric assembles the spatial frequency of Gabor texture patches and colors to simulate receptive fields, extending the subband entropy onto the object-level structure of color, orientation, and saliency [119]. This is grounded in the understanding that an image is considered ‘simple’ [98] when it is symmetric and structured. Thus, complexity is inversely correlated with perceptual grouping and regularity [51]. Empirical evaluation results also supported the FC metric (Item (5) O.FC, Table 2), subband entropy (Item (3) O.SE, Table 2) and edge density (Item (9) O.ED, Table 2) on measuring VC [119, 118]. Borgo et al. [9] further emphasized the idea of statistical attributes to characterize complexity. This line of work informs us that information amount influences VC. Both Rosenholtz in vision science and Jänicke and Chen in visualization use color as the primary channel for information-theoretic metrics [118, 65]. Given that color and textures are uniquely represented in visualization images, we modeled between continuity and complexity conditioned for continuous and discrete color and texture stimuli, or RR-ColorTexture of Continuous Heatmap and Discrete Heatmaps in images.
@@ I.2.3 Color-based Measures
Color also plays a significant role in shaping perceptual experiences. For example, Li and Chen [78] suggest that using too many colors can reduce memorability due to increased clutter. Moreover, color is a leading factor in how visual groupings are perceived, supporting both segmentation and the identification of structure within complex visual scenes. In vision science, both Greene et al. [39] and Rosenholtz [118] highlight the role of color in defining object boundaries. Many studies have linked color-related features to VC. For example, Rosenholtz et al.observed that, color variability is independent of both edge density and subband entropy [119]. Both Oliva et al. [98] and Rosenholtz et al. [120] reported that perceived VC increases more with color variability (the number of colors and how different they are) than with the presence of high frequency details in color channels. Corchs [27, 53] introduced metrics such as color harmony and color coherence to show that both the interaction and uniformity of colors contribute to VC. Here, we balance our metric choices to capture both statistical color properties and object-level color counts. Specifically, we used colorfulness (Item (7) O.CF, Table 2), defined as “the variety and diversity of colors present in the region” [44], and gray-level color heterogeneity (Item (6) O.H, Table 2) to indicate overall saturation [44], in addition to the aforementioned O.ERGB, Table 2 (see subsubsection I.2.1).
I.2.4 Object-based Measures
Several studies have focused on object-level measures, emphasizing that VC involves higher-level cognitive processing, which requires the analysis and integration of multiple elements along with their semantic content. Semantic content in visualization can have two types: graphical elements (e. g., legend, tick marks) and annotations [109] conveying key insights and takeaways [1]. The effect of texts on VC can be mixed. For one, they can effectively guide viewing behavior [13, 10] to reduce VC. For the other and from a pixel-based perspective, text is treated as a visual structure—often dark-colored, densely arranged, and high in contrast, that can increase complexity (Item (11) O.TiR, Table 2).